State administered standardized reading exam is given to eighth grade students. The scores on this exam for all students statewide have a normal distribution with a mean of 536 and a standard deviation of 70. A local Junior High principal has decided to give an award to any student who scores in the top 10% of statewide scores.

How high should a student score be to win this award?

Give your answer to the nearest integer. For help on how to input a numeric answer, please see the instructions for entering numeric response.

Please explain
Answer #1

Given that, mean (mu) = 536 and

standard deviation (sigma) = 70

We want to find, the value of x such that, P(X > x) = 0.10

P(X >x) 0.10

=> 1-P(X < x) = 0.10

> P(X < x) = 0.90

X-μ x-536 7 -) = 0.90

536 ) 0.90 70 -) = 0.90

T-536 70 = z0.90

T-536 1.28 70

> x = (1.28 * 70) 536

=> x = 89.6 + 536

=> x-025.6

=> x approx 626

Therefore, student should score 626 to win this award.